# An Adaptive Algorithm for Bilevel Optimization on Riemannian Manifolds

### v2

@@ 4.1 Simple problem
In the first experiment, following [29], we consider a simple problem, which aims to determine the maximum similarity between two matrices ${\mathbf{X}}\in{\mathbb{R}}^{n\times d}$ and ${\mathbf{Y}}\in{\mathbb{R}}^{n\times r}$, where $n\geq d\geq r$, formulated as:
@@
where ${\rm St}(d,r)=\{{\mathbf{W}}\in{\mathbb{R}}^{d\times r}:{\mathbf{W}}^{\top}{\mathbf{W}}={\mathbf{I}}_{r}\}$ and ${\mathbb{S}}_{++}^{d}=\{\mathbf{M}\in{\mathbb{R}}^{d\times d}:\mathbf{M}\succ 0\}$ represent the Stiefel manifold and the symmetric positive definite (SPD) matrices, respectively, and $\lambda>0$ is the regular parameter. The matrix ${\mathbf{W}}\in{\rm St}(d,r)$ aligns ${\mathbf{X}}$ and ${\mathbf{Y}}$ in a shared dimensional space, while the lower-level problem learns an appropriate geometric metric $\mathbf{M}\in{\mathbb{S}}_{++}^{d}$ [84]. Additionally, the geodesic strong convexity of the lower-level problem and the Hessian inverse expression can be found in Appendix H of [29]. In this experiment, we generate random data matrices ${\mathbf{X}}$ and ${\mathbf{Y}}$ with two sample sizes: $n=100$ and $n=1000$, where $d=50$ and $r=20$.
@@ 4.2 Robustness analysis
In the second experiment, also from [29, Section 4.2], we address the deep hyper-representation problem for classification, which is a subclass of hyper-representation problems [22, 63, 83, 53, 64]. In contrast to the 2-layer SPD network employed for ETH-80 dataset classification in [29], to demonstrate the efficacy of our algorithm, here we utilize a 3-layer SPD network [36] as the upper-level architecture to optimize input embeddings over the larger AFEW dataset [16], comprising seven emotion classes. The training set contains 1,747 matrices with imbalanced class distribution (267, 235, 173, 292, 342, 288, and 150 samples per class, respectively). This optimization problem is formulated as follows:
@@
where $\text{SPDnet}(\cdot;\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{A}_{3})$ denotes the 3-layer SPD network with layer parameters $\mathbf{A}_{1}$, $\mathbf{A}_{2}$, and $\mathbf{A}_{3}$ respectively [36], the term ${\mathbf{y}}_{i}$ represents the one-hot encoded label., and ${\mathcal{D}}=\{\mathbf{D}_{i}\}_{i=1}^{n}$ denote a set of SPD matrices where $\mathbf{D}_{i}\in{\mathbb{S}}_{++}^{d}$. Each matrix $\mathbf{D}_{i}$ has dimensions of $400\times 400$, and we set $d_{1}=100$, $d_{2}=20$, and $r=5$.
In this study, we perform a series of experiments with varying initial step sizes to evaluate the robustness and advantages of our proposed AdaRHD algorithm. To address computational constraints, we utilize a $5\%$ subset of the AFEW dataset [16] rather than the full dataset, ensuring tractable training durations. For AdaRHD, we initialize hyperparameters $a_{0}$, $b_{0}$, and $c_{0}$ to equal values and test performance across the range $\{0.2,1,2,10,20\}$. For RHGD, we fix $\eta_{x}$ and $\eta_{y}$ to equal values and evaluate the set $\{5,1,0.5,0.1,0.05\}$. To mitigate sampling bias and validate robustness, each algorithm is executed five times with distinct random seeds, each iteration employing a unique $5\%$ data subset. This methodology enables systematic assessment of optimization sensitivity to step sizes and establishes the generalizability of our algorithm under practical constraints.
Figure 3: Epoch vs. Validation Accuracy under different initial step sizes for each algorithm. In the figures for “ARHD-X”, the labels indicate the values of $1/a_{0}=1/b_{0}=1/c_{0}$; in the figures for “RHGD-X”, the labels represent the values of $\eta_{x}=\eta_{y}$.
Table 2: Time to Reach a Specific Validation Accuracy Under Different Initial Step Sizes for Each Algorithm. The values outside the parentheses indicate the mean over five random trials, while the values inside the parentheses represent the standard deviation. “Step Size (A/R)” denotes the initial step sizes used in each method: $1/a_{0}=1/b_{0}=1/c_{0}$ in the “ARHD-X” algorithm; $\eta_{x}=\eta_{y}$ in the “RHGD-X” algorithm. “X%” indicates the time required to reach the corresponding validation accuracy.
@@
5 280.91 468.03 1238.98 449.27 677.38 1346.98 / / / / / /
(91.02) (127.79) (669.09) (172.40) (142.76) (287.18) / / / / / /
1 128.77 187.86 241.78 213.68 320.82 425.00 / / / / / /
(19.52) (33.53) (71.50) (31.74) (55.81) (93.22) / / / / / /
0.5 115.37 167.17 237.84 192.50 293.99 466.93 / / / / / /
(18.23) (26.20) (29.97) (39.14) (59.67) (108.76) / / / / / /
0.1 222.24 319.61 436.05 306.54 428.54 541.38 165.28 358.56 789.99 33.52 158.94 527.74
(35.19) (35.02) (51.80) (55.28) (45.77) (82.30) (76.28) (70.23) (161.80) (11.87) (114.19) (62.51)
0.05 254.24 419.30 642.35 351.56 506.29 650.03 135.89 693.97 1321.19 39.41 168.10 1036.95
(41.66) (53.76) (108.29) (35.66) (25.78) (51.24) (135.61) (88.98) (175.85) (6.92) (145.02) (131.94)
@@ 5 Conclusion
This paper proposes an adaptive algorithm for solving Riemannian bilevel optimization (RBO) problems, employing two strategies: gradient descent and conjugate gradient, to compute approximate Riemannian hypergradients. To our knowledge, this is the first fully adaptive RBO algorithm, incorporating a step size mechanism eliminating prior knowledge requirements of problem parameters. We establish that the method achieves an $\mathcal{O}(1/\epsilon)$ iteration complexity to reach an $\epsilon$-stationary point, matching the complexity of the standard non-adaptive algorithms. Additionally, we show that substituting the exponential mapping with a computationally efficient retraction mapping maintains this complexity guarantee.
@@ Appendix A Related works
Bilevel optimization: Bilevel optimization is rooted in the Stackelberg game [10, 6, 75, 76]. When the lower-level objective is strongly convex, various methods have been proposed [24, 12, 38, 39, 15, 51, 32, 46]. Hypergradient computation approaches include approximate implicit differentiation (AID) [17, 58]; iterative differentiation (ITD) [54, 21, 63, 25]; Neumann series (NS) [24]; and conjugate gradient (CG) [38]. Recent studies address bilevel problems with constraints on either the upper- or lower-level objectives. For example, [13, 32] focus on scenarios where constraints are imposed solely on the upper-level objective, whereas [71, 79, 82] develop methods for problems with lower-level constraints. For further reading, readers are referred to [38, 39, 46] and references therein.
Riemannian manifold optimization: Riemannian manifold optimization has attracted considerable attention due to its broad applications, including low-rank matrix completion [8, 73]; phase retrieval [4, 67]; dictionary learning [14, 66]; dimensionality reduction [30, 70, 56]; and manifold regression [49, 50]. This methodology reformulates manifold-constrained problems into unconstrained ones by treating the manifold as the ambient space and employing retractions to compensate for the loss of linearity, thereby improving convergence properties [48]. In smooth Riemannian optimization, various methods have been developed, such as Riemannian gradient descent [23, 1, 4, 9]; nonlinear conjugate gradients [61]; and variance-reduced stochastic gradients [87, 62]. For further details, please refer to [7] and references therein.
Riemannian bilevel optimization: Research on Riemannian bilevel optimization is limited. [48] investigated hypergradient computation for such problems on Riemannian manifolds and developed deterministic and stochastic algorithms, namely RieBO and RieSBO, for solving Riemannian bilevel and stochastic bilevel optimization. In parallel, [29] proposed a framework called RHGD for Riemannian bilevel optimization, offering multiple hypergradient estimation strategies accompanied by convergence and complexity analyses. The authors also extended their framework to Riemannian minimax and compositional optimization problems. By leveraging value-function reformulation and the Lagrange method, [19] introduced a fully stochastic first-order approach applicable to scenarios where both objectives are stochastic, deterministic, or mixed.
@@ Appendix B Additional preliminaries for Section 2
[29, Definition 1] For any $x,x_{1},x_{2}\in{\mathcal{U}}_{x},y,y_{1},y_{2}\in{\mathcal{U}}_{y}$, where ${\mathcal{U}}_{x}\times{\mathcal{U}}_{y}\subseteq{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}$,
@@
Influenced by the curvature of the Riemannian manifolds, different from the distance defined in the Euclidean space, we have the following trigonometric distance bound over the Riemannian manifolds.
@@
where $\zeta(\tau,c)=\frac{\sqrt{|\tau|}c}{\tanh(\sqrt{|\tau|}c)}$ if $\tau<0$ and $\zeta(\tau,c)=1$ if $\tau\geq 0$, and $\tau$ denotes the lower bound of the sectional curvature of ${\mathcal{U}}$ [59, Section 3.1.3].
@@ Appendix C Tangent space conjugate gradient algorithm
In this section, we introduce the tangent-space conjugate gradient method [69, 7], as required in Section 3.1.
@@
2:while ${\mathcal{H}}_{y}g(x,\hat{y})[v^{n}]-{\mathcal{G}}_{y}f(x,\hat{y})>\epsilon_{v}$ do
@@
1:Initia points $x_{0}\in{\mathcal{M}}_{x},y_{0}\in{\mathcal{M}}_{y}$, and $v_{0}\in T_{y_{0}}{\mathcal{M}}$, initial step sizes $a_{0}>0$, $b_{0}>0$, and $c_{0}>0$, total iterations $T$, error tolerances $\epsilon_{y}=\epsilon_{v}=\frac{1}{T}$.
2:for $t=0,1,2,...,T-1$ do
3:     Set $p=0$ and $y_{t}^{0}=y_{t-1}^{K_{t-1}}$ if $t>0$ and $y_{0}$ otherwise.
@@
6:         $y_{t}^{k+1}={\rm Retr}_{y_{t}^{k}}(-\frac{1}{b_{k+1}}\,{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}))$,   $p=p+1.$
7:     end while
8:     $K_{t}=k$
9:     Set $n=0$ and $v_{t}^{0}={\mathcal{P}}_{y_{t-1}^{K_{t-1}}}^{y_{t}^{K_{t}}}v_{t-1}^{N_{t-1}}$ if $t>0$ and $v_{0}$ otherwise.
10:     while $\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}>{\epsilon_{v}}$ do
11:         $c_{n+1}^{2}=c_{n}^{2}+\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}$,
12:         $v_{t}^{n+1}=v_{t}^{n}-\frac{1}{c_{n+1}}\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})$$\triangleright$ Gradient descent
13:         $n=n+1$.
14:     end while
@@
16:     $N_{t}=n$,
17:     $\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})={\mathcal{G}}_{x}f(x_{t},y_{t}^{K_{t}})-{\mathcal{G}}^{2}_{xy}g(x_{t},y_{t}^{K_{t}})[v_{t}^{N_{t}}]$,
18:     $a_{t+1}^{2}=a_{t}^{2}+\|\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}$,   $x_{t+1}={\rm Retr}_{x_{t}}(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}))$.
19:end for
Since the retraction mapping only affects the total number of iterations required for the lower-level problem in Problem (1) to converge, without influencing the error bounds, the hypergradient approximation error bound (Lemma 3.1) remains consistent with that in Algorithm 1. Moreover, analogous to Proposition 3.2, we establish the following upper bounds for the total iterations required to solve the lower-level problem in Problem (1) and the linear system (4).
@@
Appendix E Extend to Riemannian mix-max problems
Riemannian min-max problems [40, 35, 28, 88, 77, 27, 55, 34] have gained increasing interest recently. The problem form of it is

### v3

@@ 4.1 Simple problem
In the first experiment, following han2024framework , we consider a simple problem, which aims to determine the maximum similarity between two matrices ${\mathbf{X}}\in{\mathbb{R}}^{n\times d}$ and ${\mathbf{Y}}\in{\mathbb{R}}^{n\times r}$, where $n\geq d\geq r$, formulated as:
@@
where ${\rm St}(d,r)=\{{\mathbf{W}}\in{\mathbb{R}}^{d\times r}:{\mathbf{W}}^{\top}{\mathbf{W}}={\mathbf{I}}_{r}\}$ and ${\mathbb{S}}_{++}^{d}=\{\mathbf{M}\in{\mathbb{R}}^{d\times d}:\mathbf{M}\succ 0\}$ represent the Stiefel manifold and the symmetric positive definite (SPD) matrices, respectively, and $\lambda>0$ is the regular parameter. The matrix ${\mathbf{W}}\in{\rm St}(d,r)$ aligns $\mathbf{X}$} and $\mathbf{Y}$} in a shared dimensional space, while the lower-level problem learns an appropriate geometric metric $\mathbf{M}\in{\mathbb{S}}_{++}^{d}$ zadeh2016geometric . Additionally, the geodesic strong convexity of the lower-level problem and the Hessian inverse expression can be found in Appendix H of han2024framework . In this experiment, we generate random data matrices $\mathbf{X}$} and $\mathbf{Y}$} with two sample sizes: $n=100$ and $n=1000$, where $d=50$ and $r=20$.
@@ 4.2 Robustness analysis
In the second experiment, also from (han2024framework, , Section 4.2), we address the deep hyper-representation problem for classification, which is a subclass of hyper-representation problems franceschi2018bilevel ; shaban2019truncated ; yu2020hyper ; lorraine2020optimizing ; sow2022convergence . In contrast to the 2-layer SPD network employed for ETH-80 dataset classification in han2024framework , to demonstrate the efficacy of our algorithm, here we utilize a 3-layer SPD network (huang2017riemannian, ) as the upper-level architecture to optimize input embeddings over the larger AFEW dataset (dhall2014emotion, ), comprising seven emotion classes. The training set contains 1,747 matrices with imbalanced class distribution (267, 235, 173, 292, 342, 288, and 150 samples per class, respectively). This optimization problem is formulated as follows:
@@
where $\text{SPDnet}(\cdot;\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{A}_{3})$ denotes the 3-layer SPD network with layer parameters $\mathbf{A}_{1}$, $\mathbf{A}_{2}$, and $\mathbf{A}_{3}$ (huang2017riemannian, ), the term ${\mathbf{y}}_{i}$ represents the one-hot encoded label, and ${\mathcal{D}}=\{\mathbf{D}_{i}\}_{i=1}^{n}$ denote a set of SPD matrices where $\mathbf{D}_{i}\in{\mathbb{S}}_{++}^{d}$. Each matrix $\mathbf{D}_{i}$ has dimensions of $400\times 400$, and we set $d_{1}=100$, $d_{2}=20$, and $r=5$.
In this study, we perform a series of experiments with varying initial step sizes to evaluate the robustness and advantages of our proposed AdaRHD algorithm. To address computational constraints, we utilize a $5\%$ subset of the AFEW dataset (dhall2014emotion, ) rather than the full dataset, ensuring tractable training durations. For AdaRHD, we initialize hyperparameters $a_{0}$, $b_{0}$, and $c_{0}$ to equal values and test performance across the range $\{0.2,1,2,10,20\}$. For RHGD, we fix $\eta_{x}$ and $\eta_{y}$ to equal values and evaluate the set $\{5,1,0.5,0.1,0.05\}$. To mitigate sampling bias and validate robustness, each algorithm is executed five times with distinct random seeds, each iteration employing a unique $5\%$ data subset. This methodology enables systematic assessment of optimization sensitivity to step sizes and establishes the generalizability of our algorithm under practical constraints.
Figure 3: Epoch vs. validation accuracy and ergodic performance $\min_{i\in[0,t]}\|\widehat{\mathcal{G}}F(x_{i},y_{i}^{K_{i}},v_{i}^{N_{i}})\|_{x_{i}}^{2}$ under different initial step sizes for each algorithm. In the figures for “AdaRHD-X”, the labels indicate the values of $1/a_{0}=1/b_{0}=1/c_{0}$; in the figures for “RHGD-X”, the labels represent the values of $\eta_{x}=\eta_{y}$.
Table 2: Time to reach a specific validation accuracy under different initial step sizes for each algorithm. The values outside the parentheses indicate the mean over five random trials, while the values inside the parentheses represent the standard deviation. “Step Size (A/R)” denotes the initial step sizes used in each method: $1/a_{0}=1/b_{0}=1/c_{0}$ in the “AdaRHD-X” algorithm; $\eta_{x}=\eta_{y}$ in the “RHGD-X” algorithm. “X%” indicates the time required to reach the corresponding validation accuracy.
@@
5.0 287.20 464.79 1217.06 432.66 690.62 1488.26 / / / / / /
$(\textbf{93.61})$ $(\textbf{124.76})$ $(\textbf{635.57})$ $(157.76)$ $(136.52)$ $(426.67)$ / / / / / /
1.0 131.84 191.51 245.63 200.26 317.31 421.99 / / / / / /
$(\textbf{18.76})$ $(\textbf{34.15})$ $(\textbf{72.42})$ $(31.21)$ $(68.92)$ $(88.31)$ / / / / / /
0.5 118.09 169.79 241.54 192.51 291.83 477.49 / / / / / /
$(\textbf{19.91})$ $(\textbf{28.59})$ $(\textbf{33.09})$ $(22.86)$ $(54.21)$ $(120.17)$ / / / / / /
0.1 217.32 312.58 427.84 297.90 447.60 546.18 171.97 370.69 808.48 36.01 168.47 540.46
$(33.96)$ $(\textbf{33.66})$ $(\textbf{53.40})$ $(42.60)$ $(53.50)$ $(66.68)$ $(78.36)$ $(68.61)$ $(146.57)$ $(\textbf{12.37})$ $(120.86)$ $(88.54)$
0.05 251.47 415.51 633.46 358.25 504.80 648.86 137.88 715.91 1510.09 41.36 179.11 1067.92
$(36.27)$ $(49.75)$ $(96.54)$ $(32.89)$ $(\textbf{15.58})$ $(\textbf{42.91})$ $(136.10)$ $(104.62)$ $(363.41)$ $(\textbf{6.88})$ $(157.76)$ $(122.98)$
@@ 5 Conclusion
This paper proposes an adaptive algorithm for solving Riemannian bilevel optimization (RBO) problems, employing two strategies: gradient descent and conjugate gradient, to compute approximate Riemannian hypergradients. To our knowledge, this is the first fully adaptive RBO algorithm, incorporating a step size mechanism, eliminating prior knowledge requirements of problem parameters. We establish that the method achieves an $\mathcal{O}(1/\epsilon)$ iteration complexity to reach an $\epsilon$-stationary point, matching the complexity of the standard non-adaptive algorithms. Additionally, we show that substituting the exponential mapping with a computationally efficient retraction mapping maintains this complexity guarantee.
Notably, this work focuses exclusively on developing an adaptive double-loop algorithm for deterministic Riemannian bilevel optimization (RBO) problems when the lower-level objective is geodesically strongly convex. Future research directions include: (1) designing single-loop adaptive algorithms (yang2024tuning, ); (2) extending the framework to stochastic settings (han2024framework, ; dutta2024riemannian, ; li2025riemannian, ) via adaptive step sizes such as inverse cumulative stochastic (hyper)gradient norms (xie2020linear, ) or other strategies (duchi2011adaptive, ; kingma2014adam, ; vaswani2019painless, ; loizou2021stochastic, ; orvieto2022dynamics, ; yun2023riemannian, ; sakai2024general, ); (3) addressing geodesically convex (non-strongly convex) lower-level objectives through regularization (alcantara2024theoretical, ) or Polyak-Łojasiewicz (PŁ) conditions (chen2024finding, ); and (4) investigating alternative adaptive step size strategies (malitsky2019adaptive, ; yun2023riemannian, ; hu2024riemannian, ; sakai2024general, ; latafat2024adaptive, ; suh2025adaptive, ) to resolve the $1/\epsilon$-gap in gradient ($G_{g}$) and Hessian-vector product ($HV_{g}$) complexities compared to non-adaptive Riemannian bilevel methods (han2024framework, ; li2025riemannian, ).
Acknowledgements
We sincerely appreciate the reviewers for their invaluable feedback and insightful suggestions. This work is partly supported by the National Key R&D Program of China under grant 2023YFA1009300, National Natural Science Foundation of China under grants 12171100, and the Major Program of NFSC (72394360,72394364).
@@ Appendix A Related works
Bilevel optimization: Bilevel optimization is rooted in the Stackelberg game [13, 8, 86, 87]. When the lower-level objective is strongly convex, various methods have been proposed [27, 16, 42, 41, 18, 56, 34, 50]. Hypergradient-based approaches include approximate implicit differentiation (AID) [20, 64]; iterative differentiation (ITD) [59, 24, 72, 28]; Neumann series (NS) [27]; and conjugate gradient (CG) [42]. Recent studies address bilevel problems with constraints on either the upper- or lower-level objective. For example, [15, 34] focus on scenarios where constraints are imposed solely on the upper-level objective, whereas [82, 90, 93] develop methods for problems with lower-level constraints. For a comprehensive overview, we refer readers to [42, 41, 50] and the references therein.
Riemannian optimization: Riemannian optimization has attracted considerable attention due to its broad applications, including low-rank matrix completion [11, 84]; phase retrieval [5, 78]; dictionary learning [17, 77]; dimensionality reduction [33, 81, 62]; and manifold regression [55, 54]. Various methods have been developed, such as Riemannian (stochastic) gradient descent [26, 1, 9, 5, 12]; nonlinear conjugate gradients [70]; and variance-reduced stochastic gradients [98, 100, 45, 71, 103]. For further details, see [10] and references therein.
Riemannian adaptive optimization: Several adaptive algorithms have been developed to solve Riemannian optimization problems. Notable examples include the Riemannian adaptive stochastic gradient algorithm (RASA) [44], Riemannian AMSGrad (RAMSGrad) [3], modified RAMSGrad [68], constrained RMSProp (cRMSProp) [67], sharpness-aware minimization on Riemannian manifolds (Riemannian SAM) [95], Riemannian natural gradient descent (RNGD) [36], and a framework for Riemannian adaptive optimization [69]. For a comprehensive overview of their application scopes, we refer readers to [69].
Riemannian bilevel optimization: Research on Riemannian bilevel optimization remains limited to a few recent studies. [53] investigated hypergradient computation for such problems on Riemannian manifolds and developed deterministic and stochastic algorithms, namely the algorithm for Riemannian bilevel optimization (RieBO) and the algorithm for Riemannian stochastic bilevel optimization (RieSBO), for solving Riemannian bilevel and stochastic bilevel optimization, respectively. In parallel, [32] proposed a framework termed Riemannian hypergradient descent (RHGD), which provides multiple hypergradient estimation strategies, supported by convergence and complexity analyses. The authors also extended their framework to address Riemannian minimax and compositional optimization problems. By leveraging the value-function reformulation and the Lagrangian method, [22] introduced a fully stochastic first-order approach, the Riemannian first-order fast stochastic approximation (RF2SA), which is applicable to scenarios where both objectives are stochastic or deterministic.
@@ Appendix B Additional preliminaries for Section 2
[32, Definition 1] For any $x,x_{1},x_{2}\in{\mathcal{U}}_{x},y,y_{1},y_{2}\in{\mathcal{U}}_{y}$, where ${\mathcal{U}}_{x}\times{\mathcal{U}}_{y}\subseteq{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}$,
@@
Due to the curvature of Riemannian manifolds, the notion of distance differs from that in Euclidean space. Accordingly, we have the following trigonometric distance bound on Riemannian manifolds.
@@
where $\zeta(\tau,c)=\frac{\sqrt{|\tau|}c}{\tanh(\sqrt{|\tau|}c)}$ if $\tau<0$ and $\zeta(\tau,c)=1$ if $\tau\geq 0$, and $\tau$ denotes the lower bound of the sectional curvature of $\mathcal{U}$} [65, Section 3.1.3].
@@ Appendix C Tangent space conjugate gradient algorithm
In this section, we introduce the tangent-space conjugate gradient method [80, 10], which is used in Section 3.1.
@@
2:while $\|{\mathcal{H}}_{y}g(x,\hat{y})[v^{n}]-{\mathcal{G}}_{y}f(x,\hat{y})\|_{\hat{y}}>\epsilon_{v}$ do
@@
2:for t=0,1,2,,T1t=0,1,2,...,T-1 do
3:  Set $k=0$ and $y_{t}^{0}=y_{t-1}^{K_{t-1}}$ if $t>0$ and $y_{0}$ otherwise.
@@
6:   $y_{t}^{k+1}={\rm Retr}_{y_{t}^{k}}(-\frac{1}{b_{k+1}}{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}))$,
7:   $k=k+1$.
8:  end while
9:  $K_{t}=k$.
13:   $v_{t}^{n+1}=v_{t}^{n}-\frac{1}{c_{n+1}}\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})$$\triangleright$ Gradient descent
14:   $n=n+1$.
15:  end while
@@
17:  $N_{t}=n$.
19:  $a_{t+1}^{2}=a_{t}^{2}+\|\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}$,
20:  $x_{t+1}={\rm Retr}_{x_{t}}(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}))$.
21:end for
Since the retraction mapping only affects the total number of iterations required for the lower-level problem in Problem (1) to converge, without influencing the error bounds, the hypergradient approximation error bound (Lemma 3.1) remains consistent with that in Algorithm 1. Moreover, analogous to Proposition 3.2, we establish the following upper bounds on the total number of iterations required to solve the lower-level problem in Problem (1) and the linear system (4).
@@
Appendix E Extension to Riemannian min-max Problems
Riemannian min-max problems [43, 38, 30, 101, 88, 29, 61, 37] have recently attracted increasing attention. The general form of such problems is